# Algebraic solution of project scheduling problems with temporal constraints

### v1

@@ 5 Solution of project scheduling problems
lemma 4.
Let $\bm{B}$ and $\bm{D}$ be matrices and $\bm{C}$ column-regular matrix such that the matrix $\bm{R}=\bm{B}\oplus\bm{D}\bm{C}$ satisfies the condition $\mathop{\mathrm{missing}}{Tr}(\bm{R})\leq\mathbb{1}$. Let $\bm{g}$ be a vector, $\bm{f}$ and $\bm{h}$ regular vectors such that the vector $\bm{s}^{-}=\bm{f}^{-}\bm{C}\oplus\bm{h}^{-}$ satisfies the condition $\bm{s}^{-}\bm{R}^{\ast}\bm{g}\leq\mathbb{1}$. Then the minimum value of the objective function in problem (23) is equal to
@@
lemma 5.
Under the same conditions as in Lemma 4, the minimum value of the objective function in problem (27) is equal to
@@ 6 Application to vaccination scheduling
The purpose of this section is to illustrate the obtained results in a real-world context of the healthcare industry. We consider problems of temporal scheduling of vaccination activities in a medical (vaccination) centre that plans to administer vaccines to its patients (see [23, 19] for further detail on vaccine administrating and scheduling).
@@
First, we note that immunization of many patients at one medical site must follow severe requirements to minimize risk for disease exposure and spread. The required measures include cleaning and disinfecting equipment, supplies and workspaces, changing and disposing gloves, needles, syringes, as well as carrying out other procedures between each patient treatment. As a result, the start times of sessions must satisfy the inequalities $b_{ij}+x_{j}\leq x_{i}$ where $b_{ij}$ is a positive parameter that represents the minimal break time required to ensure a safe transition from one vaccination to the another.
@@
A common objective of the optimal scheduling of vaccination sessions is to minimize the makespan of the vaccination plan, defined as the time between the earliest start time and the latest finish time over all sessions and written as (5). Combining with the above described constraints yields a project scheduling problem in the form of (7) or, in terms of tropical algebra, a problem at (23). A complete solution to the later problem is provided by Lemma 4.
We now present a numerical example that demonstrates computational technique involved in the solution offered by Lemma 4.
@@
The entries in the matrix $\bm{B}$, which differ from $\mathbb{0}$, set the following relationships between the start and finish times of sessions:
@@
To solve the problem by applying Lemma 4, we need to verify the conditions of the lemma. For this purpose, we evaluate the matrices
@@
$\bm{G}=\theta^{-1}(\bm{1}(\bm{1}^{T}\bm{C}\oplus\bm{1}^{T}\bm{C}\bm{R}\oplus\bm{1}^{T}\bm{C}\bm{R}^{2}\oplus\bm{R}^{3})\oplus\bm{R}\bm{1}(\bm{1}^{T}\bm{C}\oplus\bm{1}^{T}\bm{C}\bm{R}\oplus\bm{R}^{2})\\ \oplus\bm{R}^{2}\bm{1}(\bm{1}^{T}\bm{C}\oplus\bm{1}^{T}\bm{C}\bm{R})\oplus\bm{R}^{3}\bm{1}\bm{1}^{T}\bm{C})\oplus\bm{R}^{\ast}.$
@@
which provide an optimal schedule that minimize the vaccination makespan.
@@
We now assume that under the same constraints as in the previous example, the problem is to minimize the maximum absolute deviation of start times. This problem takes the form of problem (27), which is solved by Lemma 5.
@@
We note finally that one of the solutions of problem (27) coincides with the solution obtained before for problem (23). The common solution of these problems seems to be quite consistent with that as the deviation of start time decreases, the makespan of the schedule will have a general tendency to decrease as well.
@@ 7 Conclusions
This study has developed a new algebraic technique to derive direct analytical solutions for temporal project scheduling problems. We have considered a project that consists of a set of activities that are undertaken in parallel under time-based constraints (“start–start”, “start–finish” and “finish–start” precedence relations, release times, release deadlines and completion deadlines) to achieve time-based objectives (minimization of the project makespan or the maximum absolute deviation of start times of activities). Such problems occur as auxiliary tools to handle general scheduling problems that may involve non-temporal constraints (in terms of manpower, machines, tools, equipment, space, money), and are of independent interest in project management.
@@
The common approach that can be used to solve the above problems is to represent them as linear programs. The problems are then numerically solved using computational methods and techniques of linear programming (simplex method, Karmarkar algorithm), which use iterative computational procedures to obtain one of the solutions if they exist. However, this algorithmic approach does not allow to derive a direct analytical solution of the scheduling problems that describes all optimal schedules in a compact vector form.
@@
To obtain an analytical solution in this study, we have used an approach based on moving from the conventional algebra to the tropical (idempotent) algebra, which focuses on the theory and applications of algebraic systems with idempotent addition. We have started with a short self-contained introduction to tropical algebra so as to make further discussion easier to follow. We have considered a new constrained optimization problem formulated in terms of tropical algebra, and derived a complete analytical solution to the problem. Finally, the obtained result has been exploited to solve the project scheduling problems of interest, represented in the tropical algebra setting.
@@
As an outcome, new direct solutions to the scheduling problems under study have been found analytically as expressions that concisely describe all optimal schedules in parametric vector form. In contrast to the known algorithmic solutions provided by linear programming, the analytical form of the results is more insightful, handy and flexible. It provides more opportunities for formal analysis of the solutions, can tell more about the fundamental structure of the solution set and better explain the impact of the input parameters on the solutions. Moreover, the analytical representation offers a potential to refine the solution by taking into account additional constraints and objectives.
@@
Furthermore, the analytical solution obtained is more straightforward and much simpler to implement. The solution procedure consists in a finite number of matrix-vector operations and is ready for efficient software implementation to run on both serial and parallel computers. This solution is more accurate since it avoids the accumulation of rounding errors that are usually involved in numerical algorithms. At the same time, the computational complexity of the new solution, estimated as $O(n^{4})$ arithmetic operations, is quite comparable with the complexity of the Karmarkar algorithm, which is known to be $O(n^{3.5}L)$ where $L$ is the number of bits in the input [12].
@@
Real-world problems of optimal scheduling of vaccination activities in a medical centre have been described to emphasize the practical relevance and value of the research. In particular, we have shown what meaning the formal constraints and objectives of the scheduling model used can have in a real situation. Finally, simple illustrative examples have been presented in some detail to demonstrate computational technique involved in the solution obtained.

### v2

@@ 5 Solution of project scheduling problems
Theorem 4.
Let $\bm{B}$ and $\bm{D}$ be matrices and $\bm{C}$ column-regular matrix such that the matrix $\bm{R}=\bm{B}\oplus\bm{D}\bm{C}$ satisfies the condition $\mathop{\mathrm{missing}}{Tr}(\bm{R})\leq\mathbb{1}$. Let $\bm{g}$ be a vector, $\bm{f}$ and $\bm{h}$ regular vectors such that the vector $\bm{s}^{-}=\bm{f}^{-}\bm{C}\oplus\bm{h}^{-}$ satisfies the condition $\bm{s}^{-}\bm{R}^{\ast}\bm{g}\leq\mathbb{1}$. Then, the minimum value of the objective function in problem (23) is equal to
@@
Theorem 5.
Under the same conditions as in Theorem 4, the minimum value of the objective function in problem (27) is equal to
@@ 6 Application to vaccination scheduling
The purpose of this section is to illustrate the obtained results in a real-world context of the healthcare industry. We consider problems of temporal scheduling of vaccination activities in a medical (vaccination) centre that plans to administer vaccines to its patients (see [24, 20] for further detail on vaccine administration and scheduling).
@@
First, we note that immunization of many patients at one medical site must follow severe requirements to minimize risk for disease exposure and spread. The required measures include cleaning and disinfecting equipment, supplies and workspaces, changing and disposing gloves, needles, syringes, as well as carrying out other procedures between each patient treatment. As a result, the start times of sessions must satisfy the inequalities $b_{ij}+x_{j}\leq x_{i}$ where $b_{ij}$ is a positive parameter that represents the minimal break time required to ensure a safe transition from one vaccination to another.
@@
A common objective of the optimal scheduling of vaccination sessions is to minimize the makespan of the vaccination plan, defined as the time between the earliest start time and the latest finish time over all sessions and written as (5). Combining with the above described constraints yields a project scheduling problem in the form of (7) or, in terms of tropical algebra, a problem at (23). A complete solution to the later problem is provided by Theorem 4.
We now present a numerical example that demonstrates computational technique involved in the solution offered by Theorem 4.
@@
The entries in the matrix $\bm{D}$, which differ from $\mathbb{0}$, set the following relationships between the start and finish times of sessions:
@@
To solve the problem by applying Theorem 4, we need to verify the conditions of the theorem. For this purpose, we evaluate the matrices
@@
$\bm{G}=\theta^{-1}(\bm{1}\bm{1}^{T}\bm{C}\oplus\bm{1}\bm{1}^{T}\bm{C}\bm{R}\oplus\bm{1}\bm{1}^{T}\bm{C}\bm{R}^{2}\oplus\bm{1}\bm{1}^{T}\bm{C}\bm{R}^{3}\\ \oplus\bm{R}\bm{1}\bm{1}^{T}\bm{C}\oplus\bm{R}\bm{1}\bm{1}^{T}\bm{C}\bm{R}\oplus\bm{R}\bm{1}\bm{1}^{T}\bm{C}\bm{R}^{2}\\ \oplus\bm{R}^{2}\bm{1}\bm{1}^{T}\bm{C}\oplus\bm{R}^{2}\bm{1}\bm{1}^{T}\bm{C}\bm{R}\oplus\bm{R}^{3}\bm{1}\bm{1}^{T}\bm{C})\oplus\bm{R}^{\ast}.$
@@
which provide an optimal schedule that minimizes the vaccination makespan.
A graphical representation of the optimal schedule in the form of a Gantt chart is given in Figure 1.
{ganttchart}
[…]
Figure 1: The Gantt chart of the optimal schedule of $n=5$ vaccination sessions for Example 1
@@
We now assume that under the same constraints as in the previous example, the problem is to minimize the maximum absolute deviation of start times. This problem takes the form of problem (27), which is solved by Theorem 5.
@@
Note that if we take $x_{3}=4$, the obtained optimal schedule coincides with that found in the previous example. A common solution of these examples seems to be quite consistent with that as the deviation of start time decreases, the makespan of the schedule has a general tendency to decrease as well.
Figure 2 shows another solution where the optimal schedule has the start and finish times of the third session defined as $x_{3}=5$ and $y_{3}=10$.
{ganttchart}
[…]
Figure 2: The Gantt chart of the optimal schedule of $n=5$ vaccination sessions for Example 2
@@ 7 Conclusions
This study has developed a new algebraic approach to derive direct analytical solutions for temporal project scheduling problems. We have considered a project that consists of a set of activities that are undertaken in parallel under time-based constraints (“start–start”, “start–finish” and “finish–start” precedence relations, release times, release deadlines and completion deadlines) to achieve time-based objectives (minimizing the project makespan or the maximum absolute deviation of start times of activities). Such problems occur as auxiliary tools to handle general scheduling problems that may involve non-temporal constraints (in terms of manpower, machines, tools, equipment, space, money), and are of independent interest in project management.
@@
A common approach that can be used to solve the above problems is to represent them as linear programs. The problems are then numerically solved using computational methods and techniques of linear programming (simplex method, Karmarkar algorithm), which use iterative computational procedures to obtain one of the solutions if they exist. However, this algorithmic approach does not allow to derive a direct analytical solution of the scheduling problems that is capable to describe all optimal schedules in a compact vector form.
@@
To obtain an analytical solution in this study, we have used an approach based on moving from the conventional algebra to the tropical (idempotent) algebra focused on the theory and applications of algebraic systems with idempotent addition. We have started with a short self-contained introduction to tropical algebra so as to make further discussion easier to follow. We have considered a new general constrained optimization problem formulated in terms of tropical algebra, and derived a complete analytical solution to the problem. Finally, the obtained result has been exploited to solve the project scheduling problems of interest, represented in the tropical algebra setting.
@@
As an outcome, new direct solutions to the scheduling problems under study have been found analytically as expressions that concisely describe all optimal schedules in parametric vector form. In contrast to the known algorithmic solutions provided by linear programming, the analytical form of the results is more insightful, handy and flexible. It provides more opportunities for formal analysis of the solutions, can tell more about the fundamental structure of the solution set and better explain the impact of the input parameters on the solutions. Moreover, the analytical representation offers a potential to refine the solution by taking into account additional constraints and objectives, and extend the approach to multicriteria scheduling problems.
@@
Furthermore, the analytical solution obtained is more straightforward and much simpler to implement. The solution procedure consists in a finite number of matrix-vector operations and is ready for efficient software implementation to run on both serial and parallel computers. This solution is more accurate since it avoids the accumulation of rounding errors that are usually involved in numerical algorithms. At the same time, the computational complexity of the new solution, estimated as $O(n^{3})$ arithmetic operations, is less than the complexity of the Karmarkar algorithm, which is known to be $O(n^{3.5}L)$, where $L$ is the number of bits in the input [12].
@@
Real-world problems of optimal scheduling of vaccination activities in a medical centre were described to emphasize the practical relevance and value of the research. In particular, we have shown what meaning the formal constraints and objectives of the scheduling model used can have in a real situation. Finally, simple illustrative examples were presented in some detail to demonstrate computational technique involved in the solution obtained.